# Multi-Scale Protein Language Model for Unified Molecular Modeling

### v1

@@ Appendix C More Experiment Results
ms-ESM 35M 0.79 0.81 0.74 0.63 0.69 0.60
Table 9: Experimental results on molecular tasks. Compared with the vast majority of baseline models, ms-ESM performs well, which demonstrates that through the unified modeling approach we enable PLMs to perform well on pure molecule tasks as well.
@@ Appendix D More Related Work
Regarding the modality of molecules, studies on molecular modeling can be categorized into three groups. (i) 1D-based methods: These represent molecules with SMILES strings and employ language modeling techniques, such as masking and contrastive self-supervision, to enhance molecular representation (Wang et al., 2019; Honda et al., 2019; Chithrananda et al., 2020; Zhang et al., 2021a; Xue et al., 2020; Guo et al., 2022). (ii) 2D-based methods: These represent molecules with molecular graphs, sharing common ideas with general graph modeling. Some methods (Rong et al., 2020; Li et al., 2020; Zhang et al., 2021b; Li et al., 2021) mask key substructures of molecular graphs, like motifs and functional groups, and task models with reconstructing the masked parts. Others (Wang et al., 2022; Fang et al., 2022c; Lin et al., 2022a) align views from positive pairs (corrupt versions of the same graph) and simultaneously contrast views from negative pairs (different graphs). (iii) 3D-based methods: These directly utilize the 3D structure of molecules, aligning closely with our work. Earlier studies incorporated 3D information as an auxiliary input for 2D-based methods (Liu et al., 2021; Li et al., 2022b; Zhu et al., 2022; Stärk et al., 2022). More recent methods focus on molecular modeling with pure 3D inputs (Fang et al., 2022a; Zhou et al., 2023; Luo et al., 2022; Zaidi et al., 2022; Liu et al., 2022; Jiao et al., 2023). Three self-supervised techniques have been designed: geometry masking, geometry predicting, and denoising. For masking, Fang et al. (2022a) mask bond information, while Zhou et al. (2023) mask atom types, requiring models to predict masked information based on remaining context. For predicting, Fang et al. (2022a) proposes an atomic prediction task with bond information to capture global structure from local information. For denoising, models reconstruct 3D structures by adjusting corrupted structures. When corrupting structures, Zhou et al. (2023); Luo et al. (2022); Zaidi et al. (2022) add Gaussian noise to each atom of the input molecule. Several methods further introduce E(3)- and SE(3)-invariance inductive bias to the denoising technique (Zhou et al., 2023; Liu et al., 2022; Jiao et al., 2023).

### v2

@@ Appendix C More Experiment Results
ESM-AA 35M 0.79 0.81 0.74 0.63 0.69 0.60
Table 9: Experimental results on molecular tasks. Compared with the vast majority of baseline models, ESM-AA performs well, which demonstrates that through the unified modeling approach we enable PLMs to perform well on pure molecule tasks as well.
@@ Appendix D More Related Work
Regarding the modality of molecules, studies on molecular modeling can be categorized into three groups. (i) 1D-based methods: These represent molecules with SMILES strings and employ language modeling techniques, such as masking and contrastive self-supervision, to enhance molecular representation (Wang et al., 2019; Honda et al., 2019; Chithrananda et al., 2020; Zhang et al., 2021a; Xue et al., 2020; Guo et al., 2022). (ii) 2D-based methods: These represent molecules with molecular graphs, sharing common ideas with general graph modeling. Some methods (Rong et al., 2020; Li et al., 2020; Zhang et al., 2021b; Li et al., 2021; Ju et al., 2023) mask key substructures of molecular graphs, like motifs and functional groups, and task models with reconstructing the masked parts. Others (Wang et al., 2022; Fang et al., 2022c; Lin et al., 2022a) align views from positive pairs (corrupt versions of the same graph) and simultaneously contrast views from negative pairs (different graphs). (iii) 3D-based methods: These directly utilize the 3D structure of molecules, aligning closely with our work. Earlier studies incorporated 3D information as an auxiliary input for 2D-based methods (Liu et al., 2021; Li et al., 2022b; Zhu et al., 2022; Stärk et al., 2022). More recent methods focus on molecular modeling with pure 3D inputs (Fang et al., 2022a; Zhou et al., 2023; Luo et al., 2022; Zaidi et al., 2022; Liu et al., 2022; Jiao et al., 2023). Three self-supervised techniques have been designed: geometry masking, geometry predicting, and denoising. For masking, Fang et al. (2022a) mask bond information, while Zhou et al. (2023) mask atom types, requiring models to predict masked information based on remaining context. For predicting, Fang et al. (2022a) proposes an atomic prediction task with bond information to capture global structure from local information. For denoising, models reconstruct 3D structures by adjusting corrupted structures. When corrupting structures, Zhou et al. (2023); Luo et al. (2022); Zaidi et al. (2022) add Gaussian noise to each atom of the input molecule. Several methods further introduce E(3)- and SE(3)-invariance inductive bias to the denoising technique (Zhou et al., 2023; Liu et al., 2022; Jiao et al., 2023).
Appendix E Performance on the Virtual Screening Benchmarks
We conduct pre-training experiments on inter-molecule interactions and achieved strong performance in the virtual screening benchmarks. Table 10 showcases the performance of models on the DUD-E zero-shot setting. The results for the baseline methods are sourced from the DrugCLIP paper(Gao et al., 2024). As for DrugCLIP itself, we retrained it because the original DrugCLIP employed large-scale data augmentation, an operation we omitted during our retraining process. Based on the results presented in the table, we make the following observations: ESM-AA demonstrates robust performance, surpassing the majority of baseline methods, including widely used open-source virtual screening software Vina and commercial virtual screening software Glide-SP. This is due to ESM-AA’s unified modeling providing a more aligned representation space for proteins and molecules, significantly enhancing the ability to screen for high-activity molecules. Even under less-than-ideal evaluation settings, ESM-AA is only slightly surpassed by the state-of-the-art, i.e., DrugCLIP. The primary reason for this is that DrugCLIP, in addition to utilizing pocket-ligand data during its secondary pre-training, also employed a significant amount of pocket data (3.2M pockets) during its initial pre-training phase. To ensure its functionality on the DUD-E benchmark, we were unable to exclude this portion of pocket data, giving DrugCLIP an unfair advantage in comparison with ESM-AA. However, despite its inherent disadvantage, ESM-AA still achieved performance comparable to DrugCLIP, which underscores the effectiveness of its modeling strategy
E.1 Details of the Pre-training and Finetuning
Following DrugCLIP(Gao et al., 2024), we conducted secondary pre-training based on ESM-AA. This involved using protein pocket-ligand pairs as input, where pockets and ligands binding to each other served as positive samples, and randomly paired pocket-ligand combinations served as negative samples for contrastive pre-training. When processing pockets with ESM-AA, we decomposed each pocket residue into its constituent atoms, aligning with DrugCLIP’s approach. The pre-training data, comprising over 17,000 pocket-ligand complexes from PDBBind 2019, was also sourced from DrugCLIP. Hyperparameters were largely aligned with DrugCLIP, except for the learning rate, set to 1e-4 (compared to DrugCLIP’s 1e-3), as we observed that excessively high learning rates hindered ESM-AA convergence.
Consistent with DrugCLIP, we also assessed the post-secondary pre-trained ESM-AA using the challenging zero-shot setting from the DUD-E Benchmark, a widely recognized virtual screening benchmark. DUD-E encompasses 102 proteins and 22,886 bioactive molecules, each accompanied by 50 topologically dissimilar decoys with matched physicochemical properties retrieved from the ZINC database. To ensure the zero-shot setting, we excluded all targets present in the DUD-E from the pre-training set. We employed ESM-AA to extract vector representations of both pockets and ligands, leveraging cosine similarity to rank pocket-ligand pairs, with higher cosine values indicating superior ranking. Evaluation metrics included the standard area under the receiver operating characteristic curve (AUROC), Boltzmann-enhanced discrimination of the receiver operating characteristic curve (BEDROC), and Enrichment Factor (EF).
Table 10: Results on DUD-E in zero-shot setting. The details of baselines can be found in Gao et al. (2024).
Method AUROC(%) $\uparrow$ BEDROC(%) $\uparrow$ EF(0.5%$\uparrow$ EF(1%$\uparrow$ EF(5%$\uparrow$
Glide-SP 76.70 40.70 19.39 16.18 7.23
Vina 71.70 - 9.13 7.32 4.44
NN-score 68.30 12.20 4.16 4.02 3.12
RFscore 65.21 12.41 4.90 4.52 2.98
Pafnucy 63.11 16.50 4.24 3.86 3.76
OnionNet 59.71 8.62 2.84 2.84 2.20
DrugCLIP 81.72 42.24 31.12 26.23 9.83
ESM-AA 80.02 39.23 28.91 24.12 9.47
Appendix F Performance on Protein Function Annotation Tasks
We have conducted experiments on protein function annotation tasks, where ESM-AA, even without structural input, matches or exceeds the performance of structural protein representation models. Table 11 showcases the performance of models on the Protein Function Annotation Tasks.
Protein Function Annotation seeks to annotate a protein with multiple functional labels. To evaluate model performance, we leverage two established benchmarks from DeepFRI(Gligorijević et al., 2021): Enzyme Commission (EC) number prediction and Gene Ontology (GO) term prediction. The GO benchmark further categorizes predictions into three branches: molecular function (GO-MF), biological process (GO-BP), and cellular component (GO-CC). Consistent with GearNet(Zhang et al., 2022), we utilize the dataset splits with a 95% sequence identity threshold for both EC and GO predictions. Notably, all models except for those explicitly defined as structural models rely solely on protein sequences as input for all tasks, including ESM-AA .
ESM-AA demonstrates robust performance, surpassing the majority of baseline methods. Among the selected 9 baselines, ESM-AA outperforms the average performance of 8 baselines, and surpasses the ESM-2 35M model in all tasks. This demonstrates the effectiveness of our designed pretraining scheme. The ESM-AA model exhibits performance close to that of GearNet, which has the highest average performance, and outperforms the average performance of other models.
ESM-AA achieves or even surpasses the performance of structural models even without structural information input. The performance of ESM-AA surpasses that of the protein structure model (DeepFRI, New IEConv(Hermosilla & Ropinski, 2022)) and approaches the performance level of the protein structure model GearNet(Zhang et al., 2022). This indicates that even without structural information as input, ESM-AA is able to model protein semantic information effectively.
Table 11: Results on protein function annotation tasks. The details of baselines can be found in Zhang et al. (2022).
Method EC GO-BP GO-MF GO-CC
AUPR $F_{max}$ AUPR $F_{max}$ AUPR $F_{max}$ AUPR $F_{max}$
CNN 0.54 0.545 0.165 0.244 0.38 0.354 0.261 0.387
ResNet 0.137 0.187 0.166 0.28 0.281 0.267 0.266 0.403
LSTM 0.032 0.082 0.13 0.248 0.1 0.166 0.15 0.32
Transformer 0.187 0.219 0.135 0.257 0.172 0.24 0.17 0.38
ProtBert 0.859 0.838 0.188 0.279 0.464 0.456 0.234 0.408
DeepFRI 0.547 0.631 0.282 0.399 0.462 0.465 0.363 0.46
ESM-2 35M 0.803 0.786 0.274 0.384 0.582 0.584 0.32 0.395
New IEConv 0.775 0.735 0.273 0.374 0.572 0.544 0.316 0.444
GearNet 0.892 0.874 0.292 0.49 0.596 0.654 0.336 0.488
ESM-AA  35M 0.82 0.797 0.283 0.401 0.586 0.59 0.309 0.418
Appendix G More Ablation Results
Ablation on Pre-trained Model Combinations
We further analyze the performance of different protein and molecule pre-trained model combinations on the Enzyme-Substrate Affinity Regression (ESAR) task within the framework provided by ProSmith. The results are shown in Table 12. Based on the data presented in the table, we make the following observations:
  • Utilizing a unified model to process both proteins and molecules always provides better performance than using separate models to handle each independently (last row vs. other rows). Using a unified model for proteins and molecules creates more cohesive representations of both, facilitating easier alignment of corresponding protein-molecule data for downstream tasks (as illustrated in Figure 4). This approach yields better performance than employing two distinct models.
  • Even without employing ESM-AA for unified processing, using ESM-AA to handle either proteins or molecules alone can also lead to performance improvements (2nd row vs. 1st row and 3rd row vs. 1st row). We believe this is due to the implicit alignment between ESM-AA and both ESM-2 and Uni-Mol. Specifically, the loss function and training data used by ESM-AA can be considered a combination of those from ESM-2 and Uni-Mol. Furthermore, in constructing ESM-AA, we also loaded the ESM-2 checkpoint for parameter initialization. This training strategy results in an implicit alignment between ESM-AA and both ESM-2 and Uni-Mol, similarly offering an advantage in processing protein-molecule data.
Table 12: Ablation analysis of the combination of protein pre-training and molecule pre-training models. Using ESM-AA for unified protein and molecule processing yields the best performance, and performance improvements are observed even when ESM-AA is only used for proteins or molecules.
Protein Molecule ESAR
Pre-training Pre-training MSE $\downarrow$ $R^{2}\uparrow$ Pearson $\uparrow$
ESM-2 35M Uni-Mol 48M 0.642(+0.043) 0.536(-0.030) 0.733(-0.020)
ESM-AA 35M Uni-Mol 48M 0.638(+0.039) 0.539(-0.027) 0.735(-0.018)
ESM-2 35M ESM-AA 35M 0.622(+0.023) 0.550(-0.016) 0.742(-0.011)
ESM-AA 35M ESM-AA 35M 0.599 0.566 0.753